OpenStudy (anonymous):

One-third of the sum of three consecutive odd integers is −1. Find the smallest integer.

OpenStudy (anonymous):

i'm getting -7/3

OpenStudy (texaschic101):

x = 1st odd integer x + 2 = 2nd odd integer x + 4 = 3rd odd integer 1/3(x + x + 2 + x + 4) = -1 1/3(3x + 6) = -1 (divide both sides by 1/3) 3x + 6 = -1/(1/3) 3x + 6 = -1 * 3/1 3x + 6 = -3 3x = -3 - 6 3x = -9 x = -3 check... 1/3(3x + 6) = -1 1/3(3(-3) + 6) = -1 1/3(-9 + 6) = -1 1/3(- 3) = -1 -1 = -1 (correct) lowest number (x) = -3